OpenStudy (anonymous):

Can someone help me with this logarithm? Asked this earlier =( log[5](x+116)=4-log[5](x-4)

OpenStudy (vijay):

20 = (x-4) * (x+116) solve this

OpenStudy (anonymous):

are u sure its 20 and not 5^4?

OpenStudy (anonymous):

Eh.. that's wrong, sorry.

OpenStudy (anonymous):

\(log_5(x+116) = 4 - log_5(x-4)\)\[\implies log_5(x+116) + log_5(x-4) = 4\]\[\implies log_5((x+116)(x-4)) = 4\]\[\implies (x+116)(x-4) = 5^4\]

OpenStudy (vijay):

\[\log(5) (x+116) + \log(5)(x-4 ) = \log(5) (5^{4})\]

OpenStudy (anonymous):

yes polpok thats what i meant. I got to that point and im trying to solve. I got stumped. Because I know (x=116)(x-4)=625 then i need to solve and make the equation equal to zero so i can factor then correct?

OpenStudy (anonymous):

foil out the left hand side, then subtract 625 from both sides and solve the quadratic.

OpenStudy (anonymous):

According to my book i need to factor and i'm stumped at that point. i'm at.... x^2+120x-161=0 trying to factor for the final answer..lost

OpenStudy (anonymous):

If factoring seems too painful I suggest just using the quadratic formula. Do you know it?

OpenStudy (anonymous):

yes i'm just trying to go by the book.

OpenStudy (anonymous):

Although I disagree with that quadratic..

OpenStudy (anonymous):

Disagree as in it is incorrect from that point?

OpenStudy (anonymous):

You should have: \((x+116)(x-4) = 625\) \[\implies x^2 +112x - 464 = 625\]\[\implies x^2 + 112x - 1089 = 0\]

OpenStudy (anonymous):

I think you messed up your signs when you foiled

OpenStudy (anonymous):

Ugh thank you...i did miss the subtraction. i had it on paper as another plus. u caught it thanks

OpenStudy (anonymous):

@Polpak..I had a much easier time solving with the quadratic formula thank u!!